The pH of a lake is 8.0. Over a 10 year period, the pH of the water becomes 1000x more
acidic.
(a) Name an ion that has increased in the water over the 10 years? [1 Mark]
(b) What is the pH now? [3 Marks]
(c) Why does pure water have a neutral pH at 25oC? [3 Marks]
(d) The pH of pure water at 60°C is 6.51. Determine the value of Kw at this
temperature. [3 marks]
1
Expert's answer
2019-07-12T03:55:33-0400

a) H+

b) Lake pH now is 8-3=5

c) since concentration of H+ formed by water dissociation at 25C is 10^-7

d) Kw = [H+] [OH-]=[H+]²=(10^-6.51)^2=9.55*10^-14